Use the following information to answer the following questions. A study was conducted to determine if anticoagulants (warfarin) can provide greater benefit in preventing stroke than antiplatelet agents (aspirin). Of 973 patients aged ≥75 years free of stroke who were eligible and agreed to participate, 488 were randomly assigned to warfarin treatment, and 485 to aspirin treatment. They were then followed up for 3 years, during which 21 strokes occurred in warfarin group, and 44 strokes in aspirin group. (Data based on a study, 8/11/07, Lancet)

+ Stroke

-Stroke

Warfarin

21

467

488

Aspirin

44

441

485

65

908

973

Calculate the relative risk (RR) of stroke in warfarin vs. aspirin groups using warfarin as the reference group. Keep 5 decimal places for intermediate steps and round your final answer to two decimal places. (1 points)

Calculate the relative risk (RR) of stroke in warfarin vs. aspirin groups using aspirin as the reference group. Keep 5 decimal places for intermediate steps and round your final answer to two decimal places. (1 points)

In one sentence, interpret the RR you calculated from the previous question (where aspirin is the reference group). (2 points)

Calculate the risk difference between the warfarin and aspirin treatment groups using warfarin as the reference group. Keep 5 decimal places for intermediate steps and round your final answer to two decimal places. (1 points)

Calculate the risk difference between the warfarin and aspirin treatment groups using aspirin as the reference group. Keep 5 decimal places for intermediate steps and round your final answer to two decimal places. (1 points)


In one sentence, interpret the risk difference you calculated from the previous question. (2 points)
Answer #1


+Stroke

-Stroke

Warfarin

21 (A)

467 (C)

488

Aspirin

44 (B)

441 (D)

485

65

908

973

The relative risk of stroke in warfarin vs. aspirin groups using warfarin as the reference group.

RR = \left \{ A\div \left ( A+B \right ) \right \}\div \left \{ C\div \left ( C+D \right ) \right \}

= (21\div 65) \div (467\div 908)

= 0.32307\div 0.51431

RR = 0.62

(It means the risk of stroke is decreased by warfarin use.)

The relative risk of stroke in warfarin vs. aspirin groups using aspirin as the reference group.

RR = \left \{ B\div \left ( B+A \right ) \right \}\div \left \{ D\div \left ( D+C \right ) \right \}

= (44\div 65)\div (441\div 908)

= 0.67692\div 0.48568

RR = 1.39

(The risk of stroke is increased by aspirin)

risk difference between the warfarin and aspirin treatment groups using warfarin as the reference group.

RD = \left \{ A\div (A+B) \right \} - \left \{ C\div (C+D) \right \}

= (21\div 65) - (467\div 908)

= 0.32307 - 0.51431

RD = -(0.19)

Risk difference between the warfarin and aspirin treatment groups using aspirin as the reference group.

RD = \left \{ B\div (B+A) \right \} - \left \{ D\div (D+C) \right \}

= 0.67692 - 0.48568

RD = 0.19